Suppose b,c belongs to R. Define map T : R3 --> R2 by
T.(x,y,z)= (2x - 4y + 3z +b, 6xcxyz)
Show that T is linear if and only if b = c = 0.
1
Expert's answer
2020-06-11T20:35:27-0400

"T.(x,y,z)= (2x - 4y + 3z +b, 6xcxyz)"

For a mapping to be linear,

  1. "T(\alpha x,\alpha y,\alpha z)=\alpha (T(x,y,z))" "\forall\ \alpha \in \ R"
  2. "T(x_1+x_2,y_1+y_2,z_1+z_2)=T(x_1,y_1,z_1)+T(x_2,y_2,z_2)" "\forall\ \alpha \in \ R"

So, if "T" is linear,then

"T(\alpha x,\alpha y,\alpha z)=" "(2\alpha x - 4\alpha y + 3\alpha z +b, 6(\alpha x)c(\alpha x)(\alpha y)(\alpha z))=" "(2\alpha x - 4\alpha y + 3\alpha z +b," "6(\alpha^4 )cx^2yz)"

This must be equal to "(\alpha (2x - 4y + 3z +b), \alpha (6xcxyz))" .

So,"\alpha (2x - 4y + 3z +b)" "=(2\alpha x - 4\alpha y + 3\alpha z +b)"

"\implies b\alpha = b\implies b=0"

And,"\alpha (6xcxyz))" "=" "6(\alpha^4 )cx^2yz)\implies" "6\alpha x^2yzc(1-\alpha ^3)=0"

And "\alpha ,x,y,z" can have any values.

So,for the above statement to be true,"c=0."